An RC circuit have a 9V source and lies at 54seconds of the time constant. Then find the voltage across a capacitor after 153.6 seconds of charging.

1
Expert's answer
2022-04-04T17:09:29-0400
"V_C(t)=V_s\left(1-e^{-t/\tau}\right)""V_C(t)=9\:{\rm V}\left(1-e^{-153.6/54}\right)=8.48\:\rm V"